# On the integrability of extended test body dynamics around black holes

### v2

@@ Abstract
What one must do is restrict the analysis to the sub-manifold $\mathcal{T}\subset\mathcal{M}$ where the SSC $p_{\alpha}S^{\alpha\beta}=0$ holds identically. This sub-manifold has two important properties. First, it has co-dimension 3 in $\mathcal{M}$, as (3) only has 3 linearly independent components. Second, $\mathcal{T}$ is invariant under the flow of the Hamiltonian (7), as Eqs. (3) and (6) readily imply $\{C^{\alpha},H\}=-\Gamma^{\alpha}_{\beta\gamma}p^{\beta}C^{\gamma}$. Therefore, any trajectory that starts on $\mathcal{T}$ will stay on it. Both issues mentioned above (too many degrees of freedom and non-physical trajectories) are resolved by restricting our analysis to the sub-manifold $\mathcal{T}$. More precisely, we look at the intersections $\mathcal{P}:=\mathcal{T}\cap\mathcal{N}$, where the SSC holds identically and Casimir degeneracies have been lifted (this is the notion of cosymplectic sub-manifolds Zambon (2011); Vaisman (2012)). We are, therefore, in the presence of a constrained Hamiltonian system Deriglazov (2022). Expanding the SSC (3) in terms of the variables on $\mathcal{N}$ leads to $\pi_{I}D^{I}=N_{IJ}S^{I}S^{J}$ and $\pi_{0}D^{I}=\varepsilon^{I}_{\phantom{I}JK}\pi^{J}S^{K}+M_{IJ}S^{I}S^{J}$, where the matrices $N_{IJ},M_{IJ}$ only depend on $(x^{\alpha},\pi_{\alpha})$ via $\omega_{\alpha BC}$ and $\pi_{A}$. These equations readily imply $\vec{D}\cdot\vec{S}=0$ to quadratic order in spin. Geometrically, this means that $\mathcal{T}$ only intersects symplectic leaves where $\mathcal{C}_{2}=0$, and the dimension of $\mathcal{P}$ depends on the leaf $\mathcal{N}$: $\text{dim}(\mathcal{P})=10$ if $\mathcal{C}_{2}=0$ and $\text{dim}(\mathcal{P})=9$ otherwise. We focus on the former case, as it is the only one of physical interest, and let $\mathcal{N}$ be any leaf where $\mathcal{C}_{\star}=0$ from now on. It follows from the above discussion that only two out of the four SSC constraints are sufficient to define $\mathcal{P}$. In theory, any pair can be used. In practice, as in Ramond (2024), we use
$\begin{split}C^{0}&:=-\pi_{I}D^{I}+N_{IJ}S^{I}S^{J}=0\,,\\ C^{1}&:=\pi_{0}D^{1}-\varepsilon^{1}_{\phantom{I}JK}\pi^{J}S^{K}-M_{IJ}S^{I}S^{J}=0\,,\end{split}$ (13)
@@
Conserved quantities.— Our definition of integrability is the classical one –Liouville-Arnold Liouville (1855); Arnold et al. (2006)–, precisely following Hinderer and Flanagan (2008) for the relativistic context. In particular, the system $(\mathcal{P},\{,\}^{\mathcal{P}},H_{\mathcal{P}})$ will be integrable if there exists five first integrals $(\mathcal{I}_{1},\ldots,\mathcal{I}_{5})$ that are linearly independent and in pairwise involution.222 As always in a perturbative setup, the notion of integrability that we use is that of pairwise commuting first integrals up to some order in a small parameter, as used in Cárdenas-Avendaño et al. (2018); Tanay et al. (2021) and described in Fasano and Marmi (2006). A well-chosen coordinate transformation can always make this approximate notion exact Fejoz (2021). In general relativity, many spacetime of interest possess symmetries associated with Killing fields of various rank. These Killing fields are associated with conserved quantities of the MPTD system (1) and have been extensively studied in the past Dixon (1964); Rüdiger (1981, 1983); Harte (2015); Compère and Druart (2022); Harte and Dwyer (2023). We shall build upon these works to find first integrals in our Hamiltonian framework. In particular, let $k^{a}$ be a Killing vector and $Y^{ab}$ a Killing-Yano tensor of $g_{ab}$. Consider the following quantities
$\begin{split}\Xi&:=p_{a}k^{a}+\tfrac{1}{2}S^{ab}\nabla_{a}k_{b}\,,\quad\mathfrak{K}:=\tfrac{1}{4}\varepsilon_{abcd}Y^{ab}S^{cd}\,,\\ \mathfrak{Q}&:=Y^{a}_{~{}c}Y^{bc}p_{a}p_{b}+4\varepsilon_{ade[b}Y^{e}_{~{}c]}\xi^{a}S^{db}p^{c}+M_{abcd}S^{ab}S^{cd},\end{split}$ (16)
@@
where $M_{abcd}:=g_{ac}(\xi_{b}\xi_{d}-\tfrac{1}{2}g_{bd}\xi^{e}\xi_{e})-\tfrac{1}{2}Y_{a}^{\phantom{a}e}(Y_{c}^{\phantom{a}f}R_{ebfd}+\tfrac{1}{2}Y_{e}^{\phantom{e}f}R_{fbcd})$ and $\xi^{a}:=\tfrac{1}{6}\varepsilon^{abcd}\nabla_{b}Y_{cd}$. The three quantities (16) have been shown to be conserved along $\mathscr{L}$ under different assumptions about the multipolar order considered, the SSC and the background metric. We give a brief summary. First, $\Xi$ is conserved at any multipolar order, independently of any SSC, in any spacetime admitting the Killing vector $k^{a}$ Dixon (1964); Harte (2012). Second, $(\mathfrak{K},\mathfrak{Q})$ are conserved at linear order in spin, under the SSC (3), in any spacetime admitting a Killing-Yano tensor $Y^{ab}$ Rüdiger (1981, 1983); Compère and Druart (2022). Third, $(\mathfrak{K},\mathfrak{Q})$ are still conserved at quadratic order in spin, under the SSC (3), but only for a compact object with $\kappa=1$ in Eq. (2) in a Kerr background Compère et al. (2023). As the Killing invariants $(\Xi,\mathfrak{K},\mathfrak{Q})$ have been constructed at the level of the MPTD system in the literature, they are natural candidates for first integrals of the Hamiltonian system $(H_{\mathcal{P}},\{,\}^{\mathcal{P}},\mathcal{P})$.
@@
Integrability in Kerr.— We now apply the framework to the Kerr spacetime $(\mathscr{E},g_{ab})$, where $\mathscr{E}$ is covered by Boyer-Lindquist coordinates $x^{\alpha}=(t,r,\theta,\phi)$ and $g_{ab}$ is the Kerr metric with mass $M>0$ and spin parameter $a\in[0,M]$. The tetrad $(e_{A})^{a}$ that we chose is the Carter tetrad Gourgoulhon (2021); Vines et al. (2016), which admits 10 independent non-vanishing connection 1-forms $\omega_{aBC}$ Vines et al. (2016); Ramond (2024). In the Kerr spacetime, the two Killing vectors $(\partial_{t})^{a}$ and $(\partial_{\phi})^{a}$ generate the invariants $E:=-\pi_{t}$ and $L_{z}:=\pi_{\phi}$, interpreted as the object’s total energy and angular momentum component, measured at spatial infinity Gourgoulhon (2021). Similarly, the Killing-Yano tensor $Y^{ab}:=2r(e_{2})^{[a}(e_{3})^{b]}-2a\cos\theta(e_{0})^{[a}(e_{1})^{b]}$ in Kerr Floyd (1973) generates two additional invariants (16), namely $\mathfrak{K}:=r\,D^{1}+a\cos\theta\,S^{1}$, and $\mathfrak{Q}$ which is too long to be displayed here but can be found explicitly in Ramond (2024) at linear order in spin or in the Mathematica Notebook MMA (2024) at quadratic order in spin.
@@
Explicit calculations of the relevant Poisson brackets in a Kerr background revealed that: (i) $\{E,H_{\mathcal{P}}\}^{\mathcal{P}}=\{L_{z},H_{\mathcal{P}}\}^{\mathcal{P}}=0$ at linear and quadratic order in spin; (ii) $\{\mathfrak{K},H_{\mathcal{P}}\}^{\mathcal{P}}=\{\mathfrak{Q},H_{\mathcal{P}}\}^{\mathcal{P}}=0$ at linear order in spin (i.e., when setting $Q_{ABCD}=0$ in (12) and $M_{abcd}=0$ in (16)); and (iii) $\{\mathfrak{K},H_{\mathcal{P}}\}^{\mathcal{P}}$ and $\{\mathfrak{K},H_{\mathcal{P}}\}^{\mathcal{P}}$ are both proportional to $(\kappa-1)$ at quadratic order in spin. Along with $H_{\mathcal{P}}=-\tilde{\mu}^{2}/2$ itself, we thus have five first integrals $(H_{\mathcal{P}},E,L_{z},\mathfrak{K},\mathfrak{Q})$ at linear order in spin for any compact object, and at quadratic order in spin if $\kappa=1$.
@@
The five first integrals are linearly independent, and integrability will follow from their pairwise involution. To verify this, there are ten independent {,}$\{,\}^{\mathcal{P}}$-brackets to check. Nine of them involve at least one instance of either $H_{\mathcal{P}},E$ or $L_{z}$. These brackets vanish by construction (as explained in Ramond (2024)), leaving only one non-trivial bracket to compute, namely $\{\mathfrak{K},\mathfrak{Q}\}^{\mathcal{P}}$. We compute it explicitly using formulae Eqs. (15) and then applying (13). This computation, we found, is tractable by hand only to linear order in spin in a Schwarzschild background. The quadratic-in-spin case and/or Kerr cases are performed in a dedicated Mathematica Notebook MMA (2024). It reveals that $\{\mathfrak{K},\mathfrak{Q}\}^{\mathcal{P}}=0$ (even though, $\{\mathfrak{K},\mathfrak{Q}\}\neq 0$ in general, Compère and Druart (2022)). This calculation alone concludes the proof of linear-in-spin integrability in Kerr for any compact object, and of quadratic-in-spin integrability in Kerr for a compact object with $\kappa=1$. For $\kappa\neq 1$, quadratic-in-spin integrability seems compromised: There are strictly less than $5$ first integrals (we found 4 in Schwarzschild [^3] Remarkably, in Schwarzschild, $\mathfrak{Q}$ is still conserved for $\kappa\neq 1$ at quadratic-in-$S^{ab}$ order, because it coincides with the square of the total angular momentum $L^{2}$ built from the Killing vectors of the SO(3) symmetry Ramond (2024), even though the definitions of $L^{2}$ and $\mathfrak{Q}$ are completely unrelated in general. and 3 in Kerr). One may attribute this feature to objects with $\kappa\neq 1$ being somehow “less symmetric” than those with $\kappa=1$, the latter including highly “symmetric” isolated black holes.
@@
Applications to waveform modelling.— Perhaps the most crucial implication of integrability is that, to describe the linear-in-spin motion of a body orbiting a Kerr black hole, there exists 5 pairs of action-angle variables $(\vartheta^{i},\mathcal{J}_{i})_{i=1\ldots 5}$ on $\mathcal{P}$ Arnold et al. (2006). The angles $\vartheta^{i}$ are cyclic (the Hamiltonian does not depend on them) and the actions $\mathcal{J}_{i}$ are first integrals, in a 1-to-1 correspondence with the set $(H_{\mathcal{P}},E,L_{z},\mathfrak{Q},\mathfrak{K})$. This has a strong potential to improve current gravitational templates waveform generation schemes Hinderer and Flanagan (2008); Pound and Wardell (2021) developed for compact binary systems with asymmetric mass ratio, the prime targets of the LISA mission Afshordi et al. (2023).
@@
Similarly, the derivation of flux-balance laws Isoyama et al. (2019) for the integrals of motion (particularly with the recent improved understanding of hidden Kerr symmetries Grant and Moxon (2023)) and of the first law of binary mechanics in Kerr Isoyama et al. (2014); Le Tiec (2015), both relying entirely on the integrability of Kerr geodesics, can now be extended to the spinning, extended-body case. The first law (resp. flux-balance law) is critical to extract the conservative (resp. dissipative) effects of the gravitational self-force Pound et al. (2020) and benchmark the predictions of different approximation schemes Le Tiec (2014). Still regarding the self-force, our covariant Hamiltonian framework could now be used as a basis to add its conservative piece by adapting the methods of Isoyama et al. (2019); Blanco and Flanagan (2023a), and the dissipative part by using Galley’s theory of non-conservative systems Galley (2013), and investigate the integrability Nasipak (2022); Blanco and Flanagan (2023b).
@@
Future prospects.— Our Hamiltonian framework offers natural, straightforward extensions to higher multipolar orders Marsat (2015) and other quadrupolar models Steinhoff and Puetzfeld (2010); Ramond and Le Tiec (2021), in other background spacetimes. It contributes to the ongoing effort in elucidating the links between integrability and local symmetries at quadrupolar and/or quadratic-in-spin orders Frolov et al. (2017); Tanay et al. (2021); Compère et al. (2023); Harte and Dwyer (2023). The overlap of all these works would be worth investigating to reveal, perhaps, a bigger picture regarding integrability of compact objects binaries.
@@
Our result on integrability opens the possibility to extend well-known geodesic features of binary mechanics to the spinning-case. Integrability implies the existence of gauge-independent frequencies $\Omega_{i}:=\partial H/\partial\mathcal{J}_{i}$ to extend the current understanding of resonant motion of Kerr geodesics Brink et al. (2015) to spinning trajectories, now characterized by a three-fold resonances involving a spin-precession frequency in addition to azimuthal and radial components Mukherjee and Tripathy (2020). Whether quadratic-in-spin effects enhance or soften transient resonances is important for LISA, as virtually all its asymmetric binary sources will pass at least through one once Ruangsri and Hughes (2014). In addition, spin-induced quadrupoles will be invaluable to probe the state of matter in compact stars using the next generation of ground-based interferometers Maggiore et al. (2020); Hall (2022). Our framework gives a simple and powerful framework to further investigate these effects.
@@
Acknowledgments. I am indebted to A. Druart and S. Isoyama for independent verifications of the results. I also thank F. Blanco, L. Drummond, A. Harte, S. Hughes, T. Hinderer, A. Le Tiec, J. Mathews, L. Stein, S. Tanay and K. Van Aelst for interesting discussions and valuable feedback.

### v3

@@ Abstract
up to cubic-in-$S^{ab}$ corrections. Therefore, any trajectory that starts on $\mathcal{T}$ will stay on it at the perturbative ordered considered. Both issues mentioned above (too many degrees of freedom and non-physical trajectories) are resolved by restricting our analysis to the sub-manifold $\mathcal{T}$. More precisely, we look at the intersections $\mathcal{P}:=\mathcal{T}\cap\mathcal{N}$, where the SSC holds identically and Casimir degeneracies have been lifted (this is the notion of cosymplectic sub-manifolds Zambon (2011); Vaisman (2012)). Writing the SSC (3) in terms of the variables on $\mathcal{N}$ leads to $\pi_{I}D^{I}=N_{JK}S^{J}S^{K}$ and $\pi_{0}D^{I}=\varepsilon^{I}_{\phantom{I}JK}\pi^{J}S^{K}+M^{I}_{\phantom{I}JK}S^{J}S^{K}$, where $N_{JK},M^{I}_{\phantom{I}JK}$ only depend on $(x^{\alpha},\pi_{\alpha})$ via $\omega_{\alpha BC}$ and $\pi_{A}$. These equations readily imply $\vec{D}\cdot\vec{S}=0$ to quadratic order in spin. Geometrically, this means that $\mathcal{T}$ only intersects symplectic leaves where $\mathcal{C}_{2}=0$, and the dimension of $\mathcal{P}$ depends on the leaf $\mathcal{N}$: $\text{dim}(\mathcal{P})=10$ if $\mathcal{C}_{2}=0$ and $\text{dim}(\mathcal{P})=9$ otherwise. We focus on the former case, as it is the only one of physical interest, and let $\mathcal{N}$ be any leaf where $\mathcal{C}_{\star}=0$ from now on. It follows from the above discussion that only two out of the four SSC constraints are sufficient to define $\mathcal{P}$. In theory, any pair can be used. In practice, we use
$\begin{split}C^{0}&:=-\pi_{I}D^{I}+N_{IJ}S^{I}S^{J}=0\,,\\ C^{1}&:=\pi_{0}D^{1}-\varepsilon^{1}_{\phantom{1}IJ}\pi^{I}S^{J}-M^{1}_{\phantom{1}IJ}S^{I}S^{J}=0\,,\end{split}$ (14)
@@
Conserved quantities.— Our definition of integrability is the classical one –Liouville-Arnold Liouville (1855); Arnold et al. (2006)–, precisely following Hinderer and Flanagan (2008) for the relativistic context. In particular, the system $(\mathcal{P},\{,\}^{\mathcal{P}},H_{\mathcal{P}})$ will be integrable if there exists five first integrals $(\mathcal{I}_{1},\ldots,\mathcal{I}_{5})$ that are linearly independent and in pairwise involution.222 As always in a perturbative setup, the notion of integrability that we use is that of pairwise commuting first integrals up to some order in a small parameter, as used in Cárdenas-Avendaño et al. (2018); Tanay et al. (2021) and described in Fasano and Marmi (2006). In general relativity, many spacetime of interest possess symmetries associated with Killing fields of various rank. These Killing fields are associated with conserved quantities of the MPTD system (1) and have been extensively studied in the past Dixon (1964); Rüdiger (1981, 1983); Harte (2015); Compère and Druart (2022); Harte and Dwyer (2023). We shall build upon these works to find first integrals in our Hamiltonian framework. In particular, let $k^{a}$ be a Killing vector and $Y^{ab}$ a Killing-Yano tensor of $g_{ab}$. Consider the following quantities:
$\Xi$ $:=p_{a}k^{a}+\tfrac{1}{2}S^{ab}\nabla_{a}k_{b}\,,$ (17a)
$\mathfrak{K}$ $:=\tfrac{1}{4}\varepsilon_{abcd}Y^{ab}S^{cd}\,,$ (17b)
$\mathfrak{Q}$ $:=K^{ab}p_{a}p_{b}+4\varepsilon_{ade[b}Y^{e}_{~{}c]}\xi^{a}S^{db}p^{c}+M_{abcd}S^{ab}S^{cd}\,,$ (17c)
@@
where $K^{ab}:=Y^{a}_{~{}c}$ is a symmetric Killing-Stäckel tensor and $M_{abcd}:=g_{ac}(\xi_{b}\xi_{d}-\tfrac{1}{2}g_{bd}\xi^{e}\xi_{e})-\tfrac{1}{2}Y_{a}^{\phantom{a}e}(Y_{c}^{\phantom{a}f}R_{ebfd}+\tfrac{1}{2}Y_{e}^{\phantom{e}f}R_{fbcd})$, with $\xi^{a}$ defined [^3] Such that $\nabla_{a}Y_{bc}:=\xi^{d}\varepsilon_{dabc}$ and $\xi^{a}=(\partial_{t})^{a}$ in Kerr. as $\xi^{a}:=-\frac{1}{6}\varepsilon^{abcd}\nabla_{b}Y_{cd}$. The quantities $(\Xi,\mathfrak{K},\mathfrak{Q})$ have been shown to be conserved along $\mathscr{L}$ under different assumptions regarding (i) the spin order considered, (ii) the SSC, (ii) the background metric and (iv) the nature of the secondary body. These conditions are gathered from Dixon (1964); Rüdiger (1981, 1983); Harte (2012); Compère and Druart (2022); Compère et al. (2023) and summarized in Table 1. As $(\Xi,\mathfrak{K},\mathfrak{Q})$ have been constructed at the level of the MPTD system in the literature, they are natural candidates for first integrals of the Hamiltonian system $(H_{\mathcal{P}},\{,\}^{\mathcal{P}},\mathcal{P})$.
Constant      Order   SSC    Metric    Body
$\Xi$ all any any any
$\mathfrak{K},\mathfrak{Q}$ linear TD Ricci-flat any
$\mathfrak{K},\mathfrak{Q}$ quadratic TD Kerr $\kappa=1$
Table 1: Conditions for the conservation of the constants of motion.
@@
Integrability in Kerr.— We now apply the framework to the Kerr spacetime $(\mathscr{E},g_{ab})$, where $\mathscr{E}$ is covered by Boyer-Lindquist coordinates $x^{\alpha}=(t,r,\theta,\phi)$ and $g_{ab}$ is the Kerr metric with mass $M\geq 0$ and spin parameter $a\in[0,M]$. The tetrad $(e_{A})^{a}$ that we chose is the Carter tetrad, admitting 10 independent non-vanishing connection 1-forms $\omega_{aBC}$ Vines et al. (2016); Ramond (2024). The two Killing vectors $(\partial_{t})^{a}$ and $(\partial_{\phi})^{a}$ generate the invariants $E:=-\pi_{t}$ and $L_{z}:=\pi_{\phi}$, interpreted as the object’s total energy and angular momentum component, measured at spatial infinity Gourgoulhon (2021). The Killing-Yano tensor $Y^{ab}:=2r(e_{2})^{[a}(e_{3})^{b]}-2a\cos\theta(e_{0})^{[a}(e_{1})^{b]}$ Floyd (1973) generates two additional invariants (17), namely $\mathfrak{K}:=r\,D^{1}+a\cos\theta\,S^{1}$, and $\mathfrak{Q}$ which is too long to be displayed here but can be found explicitly in the Mathematica Notebook MMA (2024).
@@
Explicit calculations of the relevant Poisson brackets in a Kerr background reveal that: (i) $\{E,H_{\mathcal{P}}\}^{\mathcal{P}}=\{L_{z},H_{\mathcal{P}}\}^{\mathcal{P}}=0$ at linear and quadratic order in spin; (ii) $\{\mathfrak{K},H_{\mathcal{P}}\}^{\mathcal{P}}=\{\mathfrak{Q},H_{\mathcal{P}}\}^{\mathcal{P}}=0$ at linear order in spin (i.e., when setting $Q_{ABCD}=0$ in (12) and $M_{abcd}=0$ in (17)); and (iii) $\{\mathfrak{K},H_{\mathcal{P}}\}^{\mathcal{P}}$ and $\{\mathfrak{Q},H_{\mathcal{P}}\}^{\mathcal{P}}$ are both proportional to $(\kappa-1)$ at quadratic order in spin. Along with $H_{\mathcal{P}}=-\tilde{\mu}^{2}/2$ itself, we have five first integrals $(H_{\mathcal{P}},E,L_{z},\mathfrak{K},\mathfrak{Q})$ at linear order in spin for any compact object, and at quadratic order in spin for objects with $\kappa=1$.
@@
The five first integrals are easily shown to be linearly independent. Integrability will thus follow from their pairwise involution. To verify this, there are $\binom{5}{2}=10$ independent {,}$\{,\}^{\mathcal{P}}$-brackets to check. Nine of them involve at least one instance of either $H_{\mathcal{P}},E$ or $L_{z}$. These brackets vanish by general arguments about the Hamiltonian and Killing vectors, cf. Ramond (2024), leaving only one non-trivial bracket to compute, namely $\{\mathfrak{K},\mathfrak{Q}\}^{\mathcal{P}}$. We compute it explicitly using Eqs. (16) and then applying (14). This computation is tractable by hand to linear order in spin in a Schwarzschild background. The quadratic-in-spin case and/or Kerr cases are performed in a dedicated Mathematica Notebook MMA (2024). It reveals that, to quadratic order in spin,
$\{\mathfrak{K},\mathfrak{Q}\}^{\mathcal{P}}\propto(\kappa-1)\,,$ (18)
up to cubic-in-spin corrections, even though $\{\mathfrak{K},\mathfrak{Q}\}\neq 0$ in general, even to linear order in spin. This contrast reflects the importance of a rigorous Hamiltonian treatment of the SSC, i.e., using the $\mathcal{P}$-brackets. The result (18) alone concludes the proof of linear-in-spin integrability in Kerr for any compact object, and of quadratic-in-spin integrability in Kerr for a compact object with $\kappa=1$.
For $\kappa\neq 1$, quadratic-in-spin integrability is compromised: There are strictly less than $5$ first integrals (we found 4 in Schwarzschild [^4] Remarkably, in Schwarzschild, $\mathfrak{Q}$ is still conserved for $\kappa\neq 1$ at quadratic-in-$S^{ab}$ order, because it coincides with the square of the total angular momentum $L^{2}$ built from the Killing vectors of the SO(3) symmetry, even though the definitions of $L^{2}$ and $\mathfrak{Q}$ are completely unrelated in general. and 3 in Kerr). One may attribute this feature to objects with $\kappa\neq 1$ being, somehow, “less symmetric” than those with $\kappa=1$, the latter including highly “symmetric” isolated black holes.
But why $\kappa=1$ in particular? While still puzzling at a fundamental level, we looked precisely at intermediary steps of the calculations and observed that some quadratic-in-spin terms arising from the dipole-dipole interactions (sourced by the second term on the right-hand side of Eq. (4)) were exactly cancelled by the quadrupolar contributions (like the third term there) if and only if $\kappa=1$. This is somewhat expected: black holes’ are unique in that their multipoles satisfy non-trivial, exact relations. Thus, there can be combinations of multipole contributions of different orders that have a chance to cancel each other. To illustrate, an isolated Kerr black hole’s monopolar, dipolar and quadrupolar moments are given by $M,a$ and $Q=-a^{2}M$ Thorne (1980). Thus, schematically, one has that $[\text{quadrupole}]+[\text{monopole}]\times[\text{dipole}]^{2}=0$. For families of compact objects with multipole moments depending on their equation of state such relations between multipoles are not exact as for black holes but only approximate and microphysics dependent Yagi and Yunes (2013, 2016).
@@
Applications to waveform modelling.— Perhaps the most crucial implication of integrability is that, to describe the linear-in-spin motion of a body orbiting a Kerr black hole, there exists 5 pairs of action-angle variables $(\vartheta^{i},\mathcal{J}_{i})_{i=1\ldots 5}$ on $\mathcal{P}$ Arnold et al. (2006). The angles $\vartheta^{i}$ are cyclic (the Hamiltonian does not depend on them) and the actions $\mathcal{J}_{i}$ are first integrals, in a 1-to-1 correspondence with the set $(H_{\mathcal{P}},E,L_{z},\mathfrak{Q},\mathfrak{K})$. This has a strong potential to improve current gravitational templates waveform generation schemes Hinderer and Flanagan (2008); Pound and Wardell (2021) developed for compact binary systems with asymmetric mass ratio, a prime targets of the LISA mission Afshordi et al. (2023).
@@
Similarly, the derivation of flux-balance laws Isoyama et al. (2019) for the integrals of motion (particularly with the recent improved understanding of hidden Kerr symmetries Grant and Moxon (2023)) and of the first law of binary mechanics in Kerr Isoyama et al. (2014); Le Tiec (2015), both relying entirely on the integrability of Kerr geodesics, can now be extended to the spinning, extended-body case. These laws are critical to extract conservative and dissipative effects of the gravitational self-force Pound et al. (2020) and benchmark the predictions of different approximation schemes Le Tiec (2014). Still regarding the self-force, our covariant Hamiltonian framework could now be used as a basis to add its conservative piece by adapting the methods of Isoyama et al. (2019); Blanco and Flanagan (2023a), and the dissipative part by using Galley’s theory of non-conservative systems Galley (2013), and investigate the integrability Nasipak (2022); Blanco and Flanagan (2023b).
@@
Future prospects.— Our Hamiltonian framework offers natural, straightforward extensions, from spin effects at higher multipolar orders Marsat (2015) to tidal quadrupolar models Steinhoff and Puetzfeld (2010); Ramond and Le Tiec (2021), all in any background spacetimes. This will contribute to the ongoing effort in elucidating the links between integrability and local symmetries at quadrupolar and/or quadratic-in-spin orders Frolov et al. (2017); Tanay et al. (2021); Compère et al. (2023); Harte and Dwyer (2023). The overlap of all these works would be worth investigating to reveal, perhaps, a bigger picture regarding integrability of compact objects binaries. In addition, Hamiltonian systems are tailor-made for structure-preserving numerical schemes, such as symplectic integrators Zhong et al. (2010); Seyrich and Lukes-Gerakopoulos (2012); Wu et al. (2021). As these can substantively increase computational efficiency and long-term stability of numerical solutions, combining our framework with these numerical schemes gives a promising opportunity to investigate the quadratic-in-spin dynamics.
@@
Our result on integrability opens the possibility to extend well-known Kerr geodesic features to the spinning-case. Integrability implies the existence of gauge-independent frequencies $\Omega_{i}:=\partial H/\partial\mathcal{J}_{i}$, allowing one to extend the current understanding of resonant motion of Kerr geodesics Brink et al. (2015) to spinning trajectories, now characterized by a three-fold resonances involving a spin-precession frequency in addition to azimuthal and radial components Mukherjee and Tripathy (2020). Whether quadratic-in-spin effects enhance or soften transient resonances is important for LISA, as virtually all its asymmetric binary sources will pass at least through one once Ruangsri and Hughes (2014). In addition, spin-induced quadrupoles will be invaluable to probe the state of matter in compact stars using the next generation of ground-based interferometers Maggiore et al. (2020); Hall (2022). Our framework gives a simple and powerful framework to further investigate these effects.
@@
Acknowledgments.—I am indebted to A. Druart and S. Isoyama for independent verifications of the results. I also thank F. Blanco, L. Drummond, A. Harte, S. Hughes, T. Hinderer, A. Le Tiec, J. Mathews, L. Stein, S. Tanay and K. Van Aelst for illuminating discussions and valuable feedback.